# CogLM: Tracking Cognitive Development of Large Language Models

### v1

@@ Abstract
Piaget’s Theory of Cognitive Development (PTC) posits that the development of cognitive levels forms the foundation for human learning across various abilities. As Large Language Models (LLMs) have recently shown remarkable abilities across a wide variety of tasks, we are curious about the cognitive levels of current LLMs: to what extent they have developed and how this development has been achieved. To this end, we construct a benchmark CogLM (Cognitive Ability Evaluation for Language Model) based on PTC to assess the cognitive levels of LLMs. CogLM comprises 1,220 questions spanning 10 cognitive abilities crafted by more than 20 human experts, providing a comprehensive testbed for the cognitive levels of LLMs. Through extensive experiments across multiple mainstream LLMs with CogLM, we find that: (1) Human-like cognitive abilities have emerged in advanced LLMs (GPT-4), comparable to those of a 20-year-old human. (2) The parameter size and optimization objective are two key factors affecting the cognitive levels of LLMs. (3) The performance on downstream tasks is positively correlated with the level of cognitive abilities. These findings fill the gap in research on the cognitive abilities of LLMs, tracing the development of LLMs from a cognitive perspective and guiding the future direction of their evolution.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved impressive performance on a wide variety of Natural Language Processing (NLP) tasks, including text comprehension (Kenton and Toutanova, 2019), reasoning (Talmor et al., 2020; Webb et al., 2023), code generation (Chen et al., 2021), and mathematical problems (Fu et al., 2023). However, few studies have explored the reasons behind the evolutionary relationship among various abilities, which makes it difficult for to understand the development of LLMs’ capabilities as a whole and may pose potential risks to their further development.
@@
To explore the above questions, we construct a benchmark based on the scenario experiments used in PTC for evaluating the cognitive abilities of LLMs, denoted as CogLM. A large-scale human trial was conducted involving 207 participants aged between 6 and 20 years to ensure the alignment between the CogLM and PTC. We then perform extensive experiments on CogLM over several series of language models, including OPT (Zhang et al., 2022), Llama-2 (Touvron et al., 2023), GPT-3.5-Turbo and GPT-4 (OpenAI, 2023). Our results indicate that: (1) Advanced LLMs, such as GPT-4, have developed human-like cognitive abilities, matching those of a 20-year-old individual. (2) Two primary factors influencing these cognitive capacities in LLMs are the size of their parameters and their optimization objectives. (3) There is a positive correlation between the cognitive level of LLMs and their performance in downstream tasks.
@@ 2.1 Definition of Cognitive Abilities
Definition: Objects exist both within and outside the filed of vision and maintain a certain level of stability.
@@
Example: Q: If a plane land at 10 AM and fly for 6 hours, what time will it take off? Ans: 4 AM
@@
Example: Q: If a stone is cut into two, what will be their total mass? Ans: Not change
@@
Example: Q: Alex is excited, Paul is sad, Mike is crying, Anna is angry. The sad one is dog, the angry one is swan, the crying one is cat, the excited one is tiger. Swan likes cat, cat likes tiger, tiger likes dog, dog likes swan. What does Anna like? Ans: Cat
@@ 2.3 Consistency with Theory
After the dataset construction is completed, we consider conducting human tests to further ascertain whether CogLM is consistent with PTC and whether it can effectively reflect cognitive abilities. We randomly select 10 samples from each subset of CogLM to create questionnaires, which are then distributed to testers aged between 6 and 20. Out of the 207 completed questionnaires, 141 are deemed valid (based on the reasonableness of test duration [^1] We deem the papers completed by the questionnaires in less than 10 minutes as invalid.). We calculate the Spearman and Pearson correlation coefficients between the age of the participants and their questionnaire scores. It turns out that spearman correlation is 0.7169 and pearson correlation is 0.7362 (all the p-values $<1e-10$), indicating a strong correlation between them. This statistical result validates the effectiveness of the Standardized Annotation Guidelines we have developed in ensuring the efficacy of CogLM for assessing cognitive abilities.
@@ 3.1 Experimental Setup
We perform evaluations on the most recent and popular architectures for NLP tasks and restrict our experiments to LLMs. We conduct experiments on the popular family of GPT architecture: OPT series (Zhang et al., 2022), including models with sizes of 125M, 1.3B, 2.7B, and 6.7B, optimised for text completion; GPT-3.5-Turbo, optimised for chat; and GPT-4, whose training and architecture details are unknown (OpenAI, 2023). We also perform experiments on Llama-2 family of models(Touvron et al., 2023), including models with scale of 7B, 13B and 70B. In particular, Llama-2 series are pretrained generative language models for text completion, while Llama-2-chat is fine-tuned variation optimised for dialogue applications (see Appendix Table 7 for statistics of used LLMs). We conduct experiments on NVIDIA A100 with greedy sampling unless otherwise specified.
@@
For GPT-3.5-Turbo and GPT-4, we use the Open AI API [^2] We use ”2023-03-15-preview” version for both. to run all the evaluations. For OPT, Llama-2 and Llama-2-chat series models, we use the weights provided on the Huggingface hub [^3] https://huggingface.co/. Llama-2-chat models are used as chat-completion models, while the others are used as text-completion models. For text-completion models, as they lack the ability to follow instructions and their output format is difficult to control, we concatenate each option with the corresponding question as input, and take the option with the highest generation probability as the model’s prediction. For chat-completion models, we constrain the format of the model’s generated answers through instructions.  [^4] We set the valid output format as: “The answer is $𝑜𝑝𝑡𝑖𝑜𝑛$” in the prompt. We consider a model to provide a valid answer even if the format is incorrect. Unless specified otherwise, we always ask the model to provide a single answer with explanations.
@@ 4 Related Work
Due to the importance of LLMs, their abilities have been thoroughly evaluated on a wide range of problems. Large-scale efforts have been invested in constructing large benchmarks itegrated with numerous LM evaluations across a number of fields (Srivastava et al., 2022; Liang et al., 2022; Hendrycks et al., 2020). Due to the extremely superior performance of LLMs in a number of traditional NLP tasks, recently challenging tasks have been proposed to test the upper bound performance of LLMs (Hendrycks et al., 2021; Valmeekam et al., 2022; Gendron et al., 2023). Some benchmarks include evaluation of specific cognitive abilities, such as common sense reasoning (Ismayilzada et al., 2023), planning (Xie et al., 2024), and deductive reasoning (Saparov and He, 2022). While previous benchmarks focus on measuring either a type or a category of advanced ability in LLMs, few studies explore the development relationship between different abilities, which is crucial for understanding the emergence of LLMs’ abilities.
@@
Several works introduce tools from cognitive psychology to study LLMs. Such as understanding the behavior in LLMs (Ritter et al., 2017; Kosoy et al., 2022; Hagendorff et al., 2022), exploring the human-like abilities in LLMs (Han et al., 2022; Kosinski, 2023; Aher et al., 2023; Pan and Zeng, 2023), and improving LLMs’ performance on certain task (Betz et al., 2021). Our work is most similar to present work on using cognitive psychology to explore whether LMs “learn and think like people” by Binz and Schulz (2023), which suggests that LLMs struggle to reason causally due to the differences in how humans and LLMs learn about the world. The key difference in our approaches is that Binz and Schulz (2023) aims to study GPT-3 by assessing its advanced abilities (e.g. decision-making, information search, deliberation, and causal reasoning), while we evaluate the cognitive abilities of multiple LLMs at different stages to track the evolution process of their cognitive abilities.
@@
Theory of Cognitive Development (PTC) is the most authoritative theory in the development of psychology, developed by Jean Piaget (Piaget et al., 1952). PTC suggests that intelligence grows and develops through a series of stages. As childs interact with the world around them, they continually add new knowledge, build upon existing knowledge, and adapt previously held ideas to accommodate new information. PTC is widely used in education, psychology, linguistics, and neuroscience, providing a theoretical framework and methodology for research in these areas.
@@ 5 Conclusions
In this paper, we introduce Piaget’s Theory of Cognitive Development (PTC) as a tool to track the development of cognitive abilities of LLMs. We construct CogLM based on the scenerio experiments used in PTC, and conduct thorough human testing to ensure the alignment between CogLM and PTC. Through extensive experiments on multiple series of LLMs, we show that: (1) Human-like cognitive abilities have emerged in State-of-the-art LLMs (GPT-4), comparable to those of 20-year-old humans. (2) The parameter size and optimization objective are two key factors affecting the cognitive abilities of LLMs. (3) The ability of performing downstream tasks is positively correlated with the level of cognitive abilities. We believe that our findings can provide a novel insight into the emergence of abilities in LLMs, and shed light on the future development advanced abilities of LLMs.
@@ Appendix A Appendix
Type Series Size
Text completion OPT 125M, 1.3B, 2.7B, 6.7B
Llama-2 7B,13B,70B
Chat completion Llama-2-chat 7B,13B,70B
GPT-3.5-Turbo N/A
GPT-4 N/A
Table 7: The statistics of considered language models.

### v3

@@ Abstract
Piaget’s Theory of Cognitive Development (PTC) posits that the development of cognitive levels forms the foundation for human learning across various abilities. As Large Language Models (LLMs) have recently shown remarkable abilities across a wide variety of tasks, we are curious about the cognitive levels of current LLMs: to what extent they have developed and how this development has been achieved. To this end, we construct a benchmark CogLM (Cognitive Ability Evaluation for Language Model) based on PTC to assess the cognitive levels of LLMs. CogLM comprises 1,220 questions spanning 10 cognitive abilities crafted by more than 20 human experts, providing a comprehensive testbed for the cognitive levels of LLMs. Through extensive experiments across multiple mainstream LLMs with CogLM, we find that: (1) In our testing framework, advanced LLMs (such as GPT-4) have demonstrated human-like cognitive abilities, comparable to those of a 20-year-old human. (2) The parameter size and optimization objective are two key factors affecting the cognitive levels of LLMs. (3) The performance on downstream tasks is positively correlated with the level of cognitive abilities. These findings fill the gap in research on the cognitive abilities of LLMs, tracing the development of LLMs from a cognitive perspective and guiding the future direction of their evolution. [^1] Our code and data have been released on https://github.com/WangXinglin/CogLM.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved impressive performance on a wide variety of Natural Language Processing (NLP) tasks, including text comprehension (Kenton and Toutanova, 2019), reasoning (Talmor et al., 2020; Webb et al., 2023), code generation (Chen et al., 2021), and mathematical problems (Fu et al., 2023). However, few studies have explored the reasons behind the evolutionary relationship among various abilities, which makes it difficult to understand the development of LLMs’ capabilities as a whole and may pose potential risks to their further development.
@@
To explore the above questions, we construct a benchmark based on the scenario experiments used in PTC for evaluating the cognitive abilities of LLMs, denoted as CogLM. A large-scale human trial was conducted involving 207 participants aged between 6 and 20 years to ensure the alignment between the CogLM and PTC. We then perform extensive experiments on CogLM over several series of language models, including OPT (Zhang et al., 2022), Llama-2 (Touvron et al., 2023), GPT-3.5-Turbo and GPT-4 (OpenAI, 2023). Our results indicate that: (1) Under our testing framework, Advanced LLMs, such as GPT-4, have developed human-like cognitive abilities, matching those of a 20-year-old individual. (2) Two primary factors influencing these cognitive capacities in LLMs are the size of their parameters and their optimization objectives. (3) There is a positive correlation between the cognitive level of LLMs and their performance in downstream tasks.
@@ 2.1 Definition of Cognitive Abilities
Definition: Objects exist both within and outside the field of vision and maintain a certain level of stability.
@@
Example: Q: If a plane lands at 10 AM and flies for 6 hours, what time will it take off? Ans: 4 AM
@@
Example: Q: If a stone is cut into two, what will be their total mass? Ans: No change
@@
Example: Q: Alex is excited, Paul is sad, Mike is crying, Anna is angry. The sad one is a dog, the angry one is swan, the crying one is cat, the excited one is tiger. Swan likes cats, cat likes tiger, tiger likes dog, dog likes swan. What does Anna like? Ans: Cat
@@ 2.3 Consistency with Theory
After the dataset construction is completed, we consider conducting human tests to further ascertain whether CogLM is consistent with PTC and whether it can effectively reflect cognitive abilities. We randomly select 10 samples from each subset of CogLM to create questionnaires, which are then distributed to testers aged between 6 and 20. Out of the 207 completed questionnaires, 141 are deemed valid (based on the reasonableness of test duration [^2] We deem the papers completed by the questionnaires in less than 10 minutes as invalid. Humans received the same multiple-choice questions to answer as LLMs.). We calculate the Spearman and Pearson correlation coefficients between the age of the participants and their questionnaire scores. It turns out that spearman correlation is 0.7169 and pearson correlation is 0.7362 (all the p-values $<1e-10$), indicating a strong correlation between them. This statistical result validates the effectiveness of the Standardized Annotation Guidelines we have developed in ensuring the efficacy of CogLM for assessing cognitive abilities.
@@
Type Series Size
Text completion OPT 125M, 1.3B, 2.7B, 6.7B
Llama-2 7B,13B,70B
Chat completion Llama-2-chat 7B,13B,70B
GPT-3.5-Turbo N/A
GPT-4 N/A
Table 3: The statistics of considered language models.
@@ 3.1 Experimental Setup
We perform evaluations on the most recent and popular architectures for NLP tasks and restrict our experiments to LLMs. We conduct experiments on the popular family of GPT architecture: OPT series (Zhang et al., 2022), including models with sizes of 125M, 1.3B, 2.7B, and 6.7B, optimised for text completion; GPT-3.5-Turbo, optimised for chat; and GPT-4, whose training and architecture details are unknown (OpenAI, 2023). We also perform experiments on Llama-2 family of models(Touvron et al., 2023), including models with scale of 7B, 13B and 70B. In particular, Llama-2 series are pretrained generative language models for text completion, while Llama-2-chat is fine-tuned variation optimised for dialogue applications (see Table 3 for statistics of used LLMs). We conduct experiments on NVIDIA A100 with greedy sampling unless otherwise specified.
@@
For GPT-3.5-Turbo and GPT-4, we use the Open AI API [^3] We use ”2023-03-15-preview” version for both. to run all the evaluations. For OPT, Llama-2 and Llama-2-chat series models, we use the weights provided on the Huggingface hub [^4] https://huggingface.co/. Llama-2-chat models are used as chat-completion models, while the others are used as text-completion models. For text-completion models, as they lack the ability to follow instructions and their output format is difficult to control, we concatenate each option with the corresponding question as input, and take the option with the highest generation probability as the model’s prediction. For chat-completion models, we constrain the format of the model’s generated answers through instructions.  [^5] We set the valid output format as: “The answer is $𝑜𝑝𝑡𝑖𝑜𝑛$” in the prompt. We consider a model to provide a valid answer even if the format is incorrect. Unless specified otherwise, we always ask the model to provide a single answer with explanations. The accuracy for a stage is calculated as a macro average of the accuracies of each part of that stage.
@@ 4 Related Work
Due to the importance of LLMs, their abilities have been thoroughly evaluated on a wide range of problems. Large-scale efforts have been invested in constructing large benchmarks itegrated with numerous LM evaluations across a number of fields (Srivastava et al., 2022; Liang et al., 2022; Hendrycks et al., 2020; Biderman et al., 2023). Due to the superior performance of LLMs in a number of traditional NLP tasks, recently challenging tasks have been proposed to test the upper bound performance of LLMs (Hendrycks et al., 2021; Valmeekam et al., 2022; Gendron et al., 2023). Some benchmarks include evaluation of specific cognitive abilities, such as common sense reasoning (Ismayilzada et al., 2023), planning (Xie et al., 2024), and deductive reasoning (Saparov and He, 2022). While previous benchmarks focus on measuring either a type or a category of advanced ability in LLMs, few studies explore the development relationship between different abilities, which is crucial for understanding the emergence of LLMs’ abilities.
@@
Several works introduce tools from cognitive psychology to study LLMs. Such as understanding the behavior in LLMs (Ritter et al., 2017; Kosoy et al., 2022; Hagendorff et al., 2022; Portelance et al., 2023), exploring the human-like abilities in LLMs (Han et al., 2022; Kosinski, 2023; Aher et al., 2023; Pan and Zeng, 2023), and improving LLMs’ performance on certain task (Betz et al., 2021). Our work is most similar to present work on using cognitive psychology to explore whether LMs “learn and think like people” by Binz and Schulz (2023), which suggests that LLMs struggle to reason causally due to the differences in how humans and LLMs learn about the world. The key difference in our approaches is that Binz and Schulz (2023) aims to study GPT-3 by assessing its advanced abilities (e.g. decision-making, information search, deliberation, and causal reasoning), while we analyze the relationships between the cognitive abilities of LLMs from the perspective of development, rather than assessing the level of a single cognitive ability of LLMs. Additionally, the other concurrent work (Shah et al., 2024) considers the developmental alignment of LLMs’ performance during training to the trajectories of children’s thinking, primarily measuring the growth trajectories of various cognitive abilities, whereas our measurement of "development" focuses more on the sequence relationships of different cognitive abilities emerging.
@@
Theory of Cognitive Development (PTC) is the most authoritative theory in the development of psychology, developed by Jean Piaget (Piaget et al., 1952). PTC suggests that intelligence grows and develops through a series of stages. As children interact with the world around them, they continually add new knowledge, build upon existing knowledge, and adapt previously held ideas to accommodate new information. PTC is widely used in education, psychology, linguistics, and neuroscience, providing a theoretical framework and methodology for research in these areas.
@@ 5 Conclusions
In this paper, we introduce Piaget’s Theory of Cognitive Development (PTC) as a tool to track the development of cognitive abilities of LLMs. We construct CogLM based on the scenerio experiments used in PTC, and conduct thorough human testing to ensure the alignment between CogLM and PTC. Through extensive experiments on multiple series of LLMs, we show that: (1) In our testing framework, Human-like cognitive abilities have emerged in advanced LLMs (such as GPT-4), comparable to those of 20-year-old humans. (2) The parameter size and optimization objective are two key factors affecting the cognitive abilities of LLMs. (3) The ability of performing downstream tasks is positively correlated with the level of cognitive abilities. We believe that our findings can provide a novel insight into the emergence of abilities in LLMs, and shed light on the future development advanced abilities of LLMs.
@@
Acknowledgements
This work is supported by the Beijing Natural Science Foundation, China (Nos. 4222037, L181010).
@@ Appendix A Appendix
Model Method const early semio empat rever conse induc deduc propo plan
OPT-125M Concat -16.0 38.7 26.3 -19.2 4.0 16.2 9.3 5.6 4.0 2.2
Option -18.1 0.0 10.3 -36.9 4.0 -18.0 -4.0 2.9 -3.5 -6.0
GPT-2 Concat -20.0 41.3 30.4 -4.7 5.3 10.8 9.3 4.5 -0.5 4.8
Option -25.0 -6.7 16.7 -38.5 5.3 -18.1 -6.7 0.8 -0.5 -6.0
Table 8: Performance Comparison of "Concat" and "Option" Testing Methods on CogLM Using GPT-2 and OPT-125M.
A.1 Testing Method of Text-completion Models
For text-completion models, as they lack the ability to follow instructions and their output format is difficult to control, we concatenate each option with the corresponding question as input, and take the option with the highest generation probability as model prediction. When calculating the generation probability, we normalized the generation length to eliminate the influence of the option length. Additionally, we compared the approach of having the model interpret the questions as multiple choice and using a letter as the concatenated answer (denoted as "option") with our existing testing method (denoted as "concat") using GPT-2 and OPT-125M. The results (Table 8) show that changing the testing method from "concat" to "option" leads to a significant decrease in the performance of the text-completion model. We suppose this is due to the text-completion model being more sensitive to factors such as position bias and model preference compared to the chat-completion model. As a result, directly concatenating the options with the question and ranking them based on probability is more suitable for testing the text-completion model.